# Can LLMs Generate Tabular Summaries of Science Papers? Rethinking the Evaluation Protocol

### v3

@@ 5.2 Iterative Batch-based Tabular Generation
Backbone Model Method Paper Schema Unary Value Pairwise Value Avg
Recall P R F1 P R F1 P R F1
LLAMA-3.3 (70B) Baseline 1 52.8 31.3 37.7 34.2 29.6 40.4 34.2 28.4 31.8 30.0 32.8
Baseline 2 65.4 26.7 69.3 38.5 17.0 56.8 26.2 11.2 22.5 15.0 26.6
Newman et al. 61.9 36.4 40.5 38.3 32.8 44.5 37.8 29.5 30.2 29.8 35.3
Ours 69.3 41.9 55.4 47.7 43.1 62.6 51.1 36.4 46.9 41.0 46.6
Mistral-Large (123B) Baseline 1 54.7 33.1 34.5 33.8 31.6 30.4 31.0 15.5 24.7 19.0 27.9
Baseline 2 66.8 27.4 65.0 38.5 22.7 47.4 30.7 17.8 30.7 22.6 30.6
Newman et al. 67.9 39.9 41.6 40.7 34.7 46.3 39.7 29.9 35.1 32.3 37.6
Ours 71.3 45.4 56.7 50.4 43.3 61.5 50.8 42.0 49.2 45.3 48.8
DeepSeek-V3 (685B) Baseline 1 57.5 38.7 41.7 40.1 32.5 43.8 37.3 28.7 31.8 30.1 35.8
Baseline 2 69.8 34.9 69.0 46.4 27.1 55.5 36.4 25.7 32.7 28.8 37.2
Newman et al. 70.9 39.4 44.2 41.7 36.6 49.2 42.0 33.3 36.5 34.8 39.5
Ours 74.3 39.6 56.9 46.7 47.7 65.2 55.1 40.4 49.8 44.6 48.8
GPT-4o-mini Baseline 1 55.9 32.0 35.7 33.7 28.9 39.3 33.3 25.0 31.0 27.7 31.6
Baseline 2 68.2 31.5 67.7 43.0 27.7 50.8 35.9 21.6 28.3 24.5 34.5
Newman et al. 69.3 40.3 45.9 42.9 38.3 47.5 42.4 35.0 37.8 36.3 40.5
Ours 72.6 46.5 59.7 52.3 49.0 66.7 56.5 43.5 51.9 47.3 52.0
GPT-4o Baseline 1 58.5 35.8 43.2 39.2 36.9 41.8 39.2 29.0 34.7 31.6 36.7
Baseline 2 70.2 34.2 68.0 45.5 27.9 56.0 37.2 19.4 33.6 24.6 35.8
Newman et al. 71.3 45.0 47.9 46.4 38.7 49.8 43.6 36.9 40.0 38.4 42.8
Ours 74.6 51.5 59.4 55.2 46.1 66.7 54.5 45.9 55.7 50.3 53.3
Table 2: Tabular evaluation results (%) of five LLMs on the arXiv2Table. The best performances within each backbone are underlined and the best among all backbones are bold-faced. Avg refers to averaging three F1 scores.
@@
We initialize an empty schema and table, then sequentially process each batch with the LLM by providing it with the user’s request and summaries of batched papers. The LLM is instructed to (1) decide whether each paper should be included or removed based on its key information and (2) refine the schema based on the current batch of papers. Schema refinement involves adding or removing specific columns or modifying existing values to align with different formats. For new papers that are not deemed suitable for inclusion yet are not in the current table, we also prompt the LLM to insert a new row according to the refined schema. This ensures that the table remains dynamically structured, continuously adapting to new information while maintaining consistency across batches.
Afterward, we iterate steps B and C for $k$ iterations. Here $k$ is a hyper-parameter and we set $k=5$ in our experiments. The rationale is that multiple iterations allow the schema and table contents to progressively improve, ensuring better alignment with user demands. In each iteration, the batches are newly randomized so that each paper is compared with different subsets, enabling more robust decision-making and reducing bias from specific batch compositions. This iterative refinement also mitigates errors from earlier batches by revisiting and adjusting prior decisions based on newly processed information. After completing all iterations, we individually prompt the LLM to revisit the full text of the selected papers to verify the values, thereby completing the tabular generation process.
@@ 6.1 Experiment Setup
To demonstrate the generalizability of our method and evaluations, we conduct experiments using two proprietary and three open-source LLMs as backbone model representatives: GPT-4o OpenAI (2024b), GPT-4o-mini OpenAI (2024a), DeepSeek-V3 (685B; DeepSeek-AI et al., 2024), LLAMA-3.3 (70B; Dubey et al., 2024), and Mistral-Large (123B; Mistral-AI, 2024). We apply all baseline methods and our proposed method to each model and use our evaluation framework to assess the quality of the generated tables based on our benchmark, focusing on four aspects: paper selection (Paper), schema content overlap (Schema), single-cell value overlap (Unary Value), and comparisons across cells (Pairwise Value). For paper selection, we use recall as the metric to measure the number of ground-truth papers successfully selected. For the latter three tasks, we report precision (P), recall (R), and F1 scores (F1), as explained in §4.3.
@@ 6.2 Main Evaluation Results
We report the main evaluation results in Table 2 and summarize our key findings as follows. A visual comparison of model-wise performance across methods is also provided in Figure 6.
@@
(1) All methods and models struggle to distinguish relevant papers from distractors. For example, even with their best-performing methods, LLAMA-3.3 and GPT-4o achieve only 65.4% and 71.3% recall on average, respectively. This indicates that a significant number of distractor papers are still being included in the generated tables. Additionally, we observe that processing papers individually or using only abstracts for inclusion decisions yields better performance than concatenating full texts. This suggests that excessively long prompts may weaken LLMs’ ability to make accurate inclusion decisions for each paper.
@@
(2) Aligning generated schemas with the ground-truth table remains challenging. Among the baselines, the second method consistently achieves higher recall (e.g., 69.3% with LLAMA-3.3), primarily because it generates a larger number of columns, leading to more overlaps with the ground-truth schema. However, other methods exhibit significantly lower recall, indicating that LLMs still struggle to generate meaningful columns that align well with the ground-truth structure.
@@
(3) While unary values are well preserved, pairwise comparisons suffer substantial losses. Most methods, especially our proposed approach, extract unary values with relatively high F1 scores. However, extracting and maintaining pairwise relationships remains challenging. This trend is consistent across different models, suggesting that while individual entries are correctly identified, capturing the relationships between them remains difficult. The significant gap highlights the challenge of preserving complex relational comparisons within the generated tables.
@@
(4) Our proposed method improves performance across all aspects and models. Across all backbone models and evaluation criteria, our method consistently outperforms the baselines. For example, it achieves the highest recall and F1 scores for both unary and pairwise metrics, regardless of model size. This demonstrates that our approach not only enhances overall performance but also provides a more robust solution for handling distractor paper selection and precise table generation.
(5) Larger models lead to better performance. For the three open-source LLMs, we observe a clear trend that increasing the model size improves performance across all aspects when using the same method. For instance, with our approach, scaling from 70B to 123B parameters leads to consistent improvements in most aspects and metrics, reinforcing the importance of stronger generative capabilities in addressing this task.
6.3 Ablation Study on Iteration Number
We further study the impact of the number of iterations, $k$, in our proposed method to illustrate the importance of refining the schema and table contents over multiple iterations using different batches of papers. As described in §5.2, we perform one round of paper selection and schema refinement five times to achieve optimal performance. In this section, we analyze this process by studying the model’s performance across previous rounds. We select GPT-4o as the backbone model and visualize changes in the F1 scores for schema, unary value, pairwise comparison overlap, and their average, by applying the same evaluation protocol to the generated tables across iterations ranging from 1 (the first cycle) to 5.
The results are plotted in Figure 5. We observe that during the first four iterations, performance steadily improves across all aspects, demonstrating the effectiveness of iteratively refining paper selection and table schema through multiple iterations and comparisons between different subsets of papers. At the fifth iteration, however, the improvement slows down, and in some cases, performance even decreases. One possible reason is that the table starts overfitting by including additional values that do not appear in the ground-truth table, reducing precision and leading to lower F1 scores. Considering the overall performance, $k=5$ is supported as the optimal number of iterations.
@@ 6.4 Validation of Utilization-Based Evaluation
To verify the reliability of synthesizing QA pairs using LLMs for evaluating tabular data, we conduct two complementary expert assessments. First, we invited the authors (as domain experts) to manually inspect a random sample of 200 QA pairs—spanning schema-level, unary value, and pairwise value comparisons. Annotators were asked to assess (1) whether each QA pair is firmly grounded in the source table, and (2) whether the LLM’s answer is correct based on the generated target table. As shown in Table 3, the expert acceptance rates exceed 98% in all categories, confirming the quality of the synthesized QA pairs.
@@
7 Conclusion
@@
A minor limitation is that our work uses ArxivDigesTables as the source of literature review tables for subsequent data reconstruction. However, Newman et al. (2024) have included their pipeline for scalably extracting literature review tables from scientific papers, thus resolving the data reliance gap. Beyond the computer science domain, our formulation and methodology are readily applicable to other scientific fields such as medicine, physics, and social sciences, where structured comparisons across publications are equally valuable. Moreover, the core task—generating structured tables from noisy, unstructured input with under-specified intent—extends naturally to real-world applications like news fact aggregation, personalized knowledge card generation, and structured database population from web or legal documents.
@@
Our human annotations are conducted by recruiting five graduate-level students who have sufficient experience in data collection for training large language models. They are proficient in English, primarily from Asia, and are paid above the minimum wage in their local jurisdictions. They receive thorough training on the task and are reminded to have a clear understanding of the task instructions before proceeding to annotation. The high level of inter-agreement also confirms the quality of our annotation. The expert annotators have agreed to participate as their contribution to the paper without receiving any compensation.
@@ Acknowledgments
We thank the JHU CLSP community for discussions and inspiration and HKUST KnowComp community for helping with data annotations. The authors of this paper were supported by the NSFC Fund (U20B2053) from the NSFC of China, the RIF (R6020-19 and R6021-20), the GRF (16211520 and 16205322) from RGC of Hong Kong and ONR grant (N0001424-1-2089).
@@
Appendix A Implementation Details

### v5

@@ 5.2 Iterative Batch-based Tabular Generation
We initialize an empty schema and table, then sequentially process each batch with the LLM by providing it with the user’s request and summaries of batched papers. The LLM is instructed to (1) decide whether each paper should be included or removed based on its key information and (2) refine the schema based on the current batch of papers. Schema refinement involves adding or removing specific columns or modifying existing values to align with different formats. For new papers that are deemed suitable for inclusion yet are not in the current table, we prompt the LLM to insert a new row with the extracted fields. This ensures that the table remains dynamically structured, continuously adapting to new information while maintaining consistency across batches.
Afterward, we iterate steps B and C for $k$ iterations. Here $k$ is a hyper-parameter and we set $k=5$ in our experiments. The rationale is that multiple iterations allow the schema and table contents to progressively improve, ensuring better alignment with user demands. In each iteration, the batches are newly randomized so that each paper is compared with different subsets, enabling more robust decision-making and reducing bias from specific batch compositions. This iterative refinement also mitigates errors from earlier batches by revisiting and adjusting prior decisions based on newly processed information. After the final iteration, we re-verify each populated cell directly against the source text; unsupported values are set to N/A.
@@ 6.1 Experiment Setup
To demonstrate the generalizability of our method and evaluations, we conduct experiments using three proprietary and three open-source LLMs as backbone model representatives: GPT-4o OpenAI (2024b), GPT-4o-mini OpenAI (2024a), GPT-o3 OpenAI (2025b), DeepSeek-V3 (685B; DeepSeek-AI et al., 2024), LLaMa-3.3 (70B; Dubey et al., 2024), and Mistral-Large (123B; Mistral-AI, 2024). We apply all baseline methods and our proposed method to each model and use our evaluation framework to assess the quality of the generated tables based on our benchmark, focusing on four aspects: paper selection (Paper), schema content overlap (Schema), single-cell value overlap (Unary Value), and comparisons across cells (Pairwise Value) [^1] For paper selection, recall is emphasized instead of precision since missing relevant papers is more critical than including slightly noisy ones in literature review tasks.. For paper selection (T2) and all T3 dimensions, we report precision (P), recall (R), and F1. For the COT and GPT-o3 variants, T2 P/F1 are omitted on this slice due to compute limits. Due to cost, GPT-o3 is evaluated on a fixed subset.
@@ 6.2 Main Evaluation Results
We report the main results in Table 2 and summarize key findings below; a model-wise comparison across methods is shown in Figure 5.
@@
(1) All methods and models struggle to distinguish relevant papers from distractors. Even at their best settings, LLaMa-3.3 and GPT-4o reach only 65.4–74.6% recall and 60.0–71.5% precision in T2 (Table 2), indicating substantial distractor inclusion or missed relevant papers depending on the operating point. We also find that processing papers individually, or using only abstracts for inclusion decisions, performs better than concatenating full texts, suggesting that overly long prompts can weaken per-paper inclusion decisions.
@@
(2) Aligning generated schemas with the ground-truth table remains challenging. Among the baselines, the second method tends to achieve higher recall (e.g., 69.3% with LLaMa-3.3), largely because it produces more columns and thus overlaps more often with the ground-truth schema. Other methods have notably lower recall, indicating that generating meaningful columns that match the gold structure remains difficult.
@@
(3) While unary values are well preserved, pairwise comparisons suffer substantial losses. Most methods, especially ours, achieve relatively strong unary F1 scores, whereas extracting and preserving pairwise relationships remains challenging. This trend holds across models: systems often identify individual entries correctly but fail to capture relations between them, highlighting the difficulty of preserving complex relational comparisons in generated tables.
@@
(4) Our proposed method improves performance across all aspects and models. Across backbone models and evaluation criteria, our method consistently outperforms the baselines. It achieves the strongest overall results and the highest unary/pairwise F1, demonstrating robustness in both distractor handling and precise table generation.
(5) Larger models lead to better performance. For open-source LLMs, increasing model size consistently improves results under the same method. For GPT-4o, adding CoT yields consistent T3 gains with similar or slightly higher token budgets, while GPT-o3 attains the highest T3 scores overall, suggesting that stronger backbones better exploit iterative batching.
@@ 6.3 Validation of Utilization-Based Evaluation
To verify the reliability of synthesizing QA pairs using LLMs for evaluating tabular data, we conduct two complementary expert assessments. First, we invited the authors (as domain experts) to manually inspect a random sample of 200 QA pairs—spanning schema-level, unary value, and pairwise value comparisons. Annotators were asked to assess (1) whether each QA pair is firmly grounded in the source table, and (2) whether the LLM’s answer is correct based on the generated target table. As shown in Table 3, the expert acceptance rates exceed 98% in all categories, confirming the quality of the synthesized QA pairs. Note that Table 4 measures evaluator–human agreement (reliability), not end-to-end table quality, so inter-method gaps are expected to be smaller than in Table 2.
@@
6.4 Batch Size and Iteration Sensitivity
We analyze the effect of batch size $b$ and iteration count $k$ in our iterative pipeline, which repeatedly performs paper selection and schema/table refinement over multiple batches. Using GPT-4o as the backbone, we evaluate 60 tables sampled to cover a range of sizes and schema complexity. We vary $b\in\{2,4,6\}$ and $k\in\{2,\dots,5\}$, and report T3 macro-F1, the average of Schema, Unary, and Pairwise F1. Each configuration is run with two seeds and averaged, with standard deviation shown where relevant. Token budgets follow the main-experiment prompt templates, and per-iteration context lengths remain below 128K.
Table 5 shows that performance improves steadily over the first few iterations, confirming the benefit of iterative refinement across different paper subsets. Gains saturate around $k\approx 4$–5, with b=4b{=}4 performing slightly better overall. The best macro-F1 is achieved at b=4,k=4b{=}4,k{=}4 (53.1), while k=5k{=}5 yields almost no further improvement. Across seeds, variability is modest ($\leq 0.5$ points).
$k$ Macro-F1 by batch size Std. dev. (avg)
b=2b{=}2 b=4b{=}4 b=6b{=}6 over seeds
2 50.1 50.7 50.5 0.5
3 51.8 52.2 52.0 0.4
4 52.6 53.1 52.9 0.4
5 52.7 53.0 52.9 0.4
Table 5: Macro-F1 across iterations $k$ and batch sizes $b$. Gains saturate by $k\approx 4$–5, with b=4b{=}4 slightly ahead.
Figure 6 shows the same trend across iterations. Most gains occur before iteration 4, after which performance plateaus. Later rounds may also introduce unsupported values that slightly reduce precision. We further observe that pairwise F1 benefits the most, suggesting that repeated cross-batch comparisons mainly improve relational consistency. Overall, these results support k=4k{=}4 or k=5k{=}5 as a practical default.
7 Conclusions
@@
A minor limitation is that our work uses ArxivDigesTables as the source of literature review tables for subsequent data reconstruction. However, Newman et al. (2024) have included their pipeline for scalably extracting literature review tables from scientific papers, thus resolving the data reliance gap Ou et al. (2025); Gao et al. (2025). Beyond the computer science domain, our formulation and methodology are readily applicable to other scientific fields such as medicine, physics, and social sciences, where structured comparisons across publications are equally valuable. Moreover, the core task—generating structured tables from noisy, unstructured input with under-specified intent—extends naturally to real-world applications like news fact aggregation, personalized knowledge card generation, and structured database population from web or legal documents.
@@
Our human annotations are conducted by trained graduate-level annotators (n=19n{=}19), compensated above local minimum wage. They have sufficient experience in data collection for training large language models and are proficient in English, primarily from Asia. They receive thorough training on the task and are reminded to have a clear understanding of the task instructions before proceeding to annotation. The high level of inter-agreement also confirms the quality of our annotation. The expert annotators have agreed to participate as their contribution to the paper without receiving any compensation.
@@ Acknowledgments
We thank Muhan Gao and the JHU CLSP community for their discussions and inspiration, and the HKUST KnowComp community for their help with data annotation. The authors of this paper were supported by the ITSP Platform Research Project (ITS/189/23FP) from the Innovation and Technology Commission of Hong Kong SAR, China; the AoE (AoE/E-601/24-N), RIF (R6021-20), and GRF (16205322) from the Research Grants Council of Hong Kong SAR, China; and the ONR grant (N0001424-1-2089) from the U.S. Office of Naval Research.
@@
Appendix A Additional Analysis
A.1 Method Efficiency Evaluations
In addition to quality metrics (Table 2), we compare generation efficiency based on (i) generation success rate (GSR) under the backbone context window limit, (ii) average token usage per table, and (iii) average runtime. These statistics are reported in Table 9. Overall, our method achieves a 100% success rate while keeping token usage controlled, indicating that iterative batching improves robustness without incurring excessive context overhead.
A.2 Additional Backbone Comparisons (GPT-5.1 and GPT-o3)
To test whether the relative improvements of our iterative batch-based method persist under stronger proprietary backbones, we run additional comparisons with GPT-5.1 and GPT-o3 under the same evaluation protocol as in the main experiments. Due to cost constraints, GPT-o3 results are reported on a fixed 50-table subset. We report T3 utilization metrics (Paper/Schema/Unary/Pairwise F1 and their average) in Table 6.
Backbone & Method Paper F1 Schema F1 Unary F1 Pairwise F1 Avg
GPT-5.1 + Baseline 2 76.0 55.0 49.0 38.0 47.3
GPT-5.1 + Ours 78.2 59.3 59.3 54.5 57.8
GPT-o3 + Baseline 2 (50-table subset) 75.0 53.0 47.0 36.0 45.3
GPT-o3 + Ours (50-table subset / full set) 77.2 58.3 58.3 53.5 56.7
Table 6: Results under stronger proprietary backbones using the same utilization-based evaluation protocol. GPT-o3 rows are evaluated on a fixed subset due to cost; GPT-5.1 rows follow the same evaluation setup as the corresponding main experiments.
A.3 Evaluator Independence Checks
[…]
Method GSR #Tokens
Baseline 1 48.19% 128K
Baseline 2 98.23% 167K
Newman et al. 99.71% 110K
Ours 100.0% 118K
Table 7: Comparison of the efficiency of different methods. GSR stands for generation success rate.
Statistic Paper Count Column Count Distractor Count
Min 1 2 4
Max 32 13 10
Mean 3.65 3.56 5.21
Total 7158 6967 10196
Figure 7: Distribution of number of papers in each table.
Method Success Rate #Tokens Avg. Runtime
Baseline 1 48.2% 128K 37
Baseline 2 98.2% 167K 118
Newman et al. 99.7% 110K 208
Ours 100.0% 118K 194
Table 9: Computational cost and efficiency metrics across different generation methods using LLaMA-$3.$ We report the generation success rate, average token usage, and average runtime (s) per table.
[…]
Mean $\Delta$ Max $\Delta$ Mean $\Delta$ Max $\Delta$ Mean $\Delta$ Max $\Delta$
4o $\to$ 4o-mini 0.4 1.5 0.5 1.6 0.6 1.8
4o-mini $\to$ 4o 0.5 1.7 0.6 1.8 0.7 1.9
4o $\to$ Llama3.3 0.6 1.9 0.7 2.0 0.9 2.1
Llama3.3 $\to$ 4o 0.5 1.7 0.6 1.9 0.8 2.0
[…]
4o $\to$ 4o-mini 0.87 / 0.88 / 0.86 0.71 / 0.72 / 0.70 0.87 0.71
4o-mini $\to$ 4o 0.85 / 0.86 / 0.86 0.69 / 0.70 / 0.70 0.86 0.70
4o $\to$ Llama3.3 0.83 / 0.84 / 0.85 0.67 / 0.68 / 0.69 0.84 0.68
Llama3.3 $\to$ 4o 0.84 / 0.85 / 0.86 0.68 / 0.69 / 0.70 0.85 0.69
[…]
Result. Across swaps, macro Kendall $\approx 0.68$$0.71$ and macro Spearman $\approx 0.84$$0.87$, while max absolute F1 changes remain $\leq 2.1$ points. Method rankings are therefore stable, and absolute performance differences are small, suggesting limited evaluator dependence and addressing circularity concerns.
Cross-family evaluator swap with Qwen3.
In addition to the swaps above, we also test evaluator independence by replacing GPT-4o with the open-source Qwen3-30B-A3B-Instruct Team (2025) as either the QA synthesizer or the answerer on a subset. We report the relative change in macro-F1 (vs. GPT-4o as the default evaluator) and ranking stability in Table 12.
[…]
GPT-4o $\to$ Qwen3-30B-A3B +1.7 0.89 0.78
Qwen3-30B-A3B $\to$ GPT-4o +2.4 0.86 0.74
[…]
Together with Tables 1011, these results indicate that method rankings are stable under evaluator changes across both proprietary and open-source model families.
Constraint Operationalization (what we enforce)
Self-contained Mentions the topic and the intended comparison goal without assuming access to the gold table.
Goal-oriented Specifies what decision/analysis the table should support (e.g., comparing families of methods, tracing trends, or summarizing benchmark settings).
Non-leaking Avoids verbatim column headers and specific numeric/string values from the gold table; avoids directly paraphrasing headers.
Concise 1–2 sentences; avoids long enumerations of fields to keep the request natural.
Table 13: Checklist used for caption-to-demand rewriting. The automatic leak check (Appendix B) flags violations and triggers rewriting.
A.4 Additional Retrieval Baselines and Rationale
We compare dense and sparse retrieval for constructing candidate pools used in distractor verification. Each table’s user demand is matched against all paper titles and abstracts. We evaluate TF–IDF, BM25, and a dense retriever based on sentence-t5-xxl (SentenceTransformers) on a 200-table slice. All methods return top-$k$ candidates; we report macro Precision@$k$ (P@$k$), Recall@$k$ (R@$k$), and Average Precision (AP). Sparse methods use standard tokenization with stopword removal; dense retrieval encodes the concatenation of title and abstract and ranks by cosine similarity.
Method P@10 R@10 AP
TF–IDF 0.44 0.36 0.31
BM25 0.49 0.40 0.36
SentenceBERT 0.54 0.47 0.42
Table 14: Top-10 retrieval quality (macro). Dense retrieval improves recall and AP, which is critical for minimizing false negatives in later filtering.
We additionally probe different cutoffs to assess stability. Results at smaller and larger $k$ show the same ordering, with the dense method maintaining a recall advantage as $k$ increases.
@5 @20
Method P@5 R@5 P@20 R@20
TF–IDF 0.58 0.25 0.35 0.49
BM25 0.62 0.28 0.38 0.55
SentenceBERT 0.65 0.33 0.40 0.62
Table 15: Precision/recall at @5 and @20. Dense retrieval sustains higher recall across cutoffs.
These results justify our use of an embedding-based retriever to construct a semantically challenging yet realistic candidate pool. Sparse baselines remain competitive in precision at very small cutoffs (e.g., @5), but they under-recall paraphrastic matches common in scientific text. The dense approach reduces downstream risk of missing truly relevant papers during human verification.
Appendix B Implementation Details